OpenStudy (anonymous):

64x^3-1=0 find the real or imaginary solutions by factoring

OpenStudy (internettroll11):

you have a 'difference of cubes' 1=1^3 64=4^3 take (x^3-y^3) = (x-y)(x^2+xy+y^2) so (64x^3-1)=(4x-1)(16x^2+4x+1)

OpenStudy (internettroll11):

the rest should be easy to do hopefully.

OpenStudy (anonymous):

Factoring: 64x3-1 To find the real root Theory : A difference of two perfect cubes, a3 - b3 can be factored into (a-b) • (a2 +ab +b2) Proof : (a-b)•(a2+ab+b2) = a3+a2b+ab2-ba2-b2a-b3 = a3+(a2b-ba2)+(ab2-b2a)-b3 = a3+0+0+b3 = a3+b3 Check : 64 is the cube of 4 Check : 1 is the cube of 1 Check : x3 is the cube of x1 Factorization is : (4x - 1) • (16x2 + 4x + 1) Solve for x, 4x-1= 0 x=1/4 only real root

OpenStudy (anonymous):

Yet we still have to find imaginary roots, yes?

OpenStudy (anonymous):

@Nexxus do you know how to find those from the info. given? Or still need help :)